# Ultrastable jammed sphere packings with a wide range of particle dispersities

### v1

@@ Abstract
We show that for a standard continuously-polydisperse model with particle-diameter distribution $P(\sigma)\propto\sigma^{-3}$ and polydispersity index $\Delta$, employing a combination of standard SWAP moves and transient degree of freedom (TDOF) moves during a Lubachevsky-Stillinger-like particle-growth process dramatically increases the generated packings’ jamming densities $\phi_{\rm J}(\Delta)$ and coordination numbers $Z_{\rm J}(\Delta)$, for a wide range of $\Delta$. We find that the fractional increase in $\phi_{\rm J}(\Delta)$ obtained by employing these moves first increases rapidly with $\Delta$, then plateaus at $6-7\%$ over the range $0.10\lesssim\Delta\leq 50$; the obtained $\phi_{\rm J}$ are as high as $0.747$ (for $\Delta=0.50$). These density increases are achieved without producing crystallization or fractionation. SWAP and TDOF moves also reduce packings’ rattler populations by as much as 96% and increase their bulk moduli by as much as 154%.
@@ I Introduction
Issues like these led Torquato et al. to ask the question “Is random close packing well-defined?”, and demonstrate using simulations that it is not Torquato et al. (2000). They found that the varying the particle-growth rate $\Gamma$ in the Lubachevsky-Stillinger (LS) algorithm Lubachevsky et al. (1991) produces monodisperse frictionless-hard-sphere (MFHS) packings with a wide range of jamming densities $0.64\lesssim\phi_{\rm J}\lesssim 0.68$, with larger $\Gamma$ producing lower $\phi_{\rm J}$. Using these results, they proposed that jammed packings exist over a finite domain in ($\phi_{\rm J},\Psi$)-space, where $\Psi$ is some (as yet undefined) scalar order parameter, and that the traditional estimate of $\phi_{\rm RCP}$ (0.637 Scott (1960)) actually corresponds to the “MRJ” packing fraction $\phi_{\rm MRJ}$, i.e, the packing fraction of maximally random jammed states that minimize $\Psi$ Torquato et al. (2000); Donev et al. (2004). In simulations of MFHS, $\phi_{\rm MRJ}$ is the jamming density obtained by starting from a state with a packing fraction below the “onset” density $\phi_{\rm on}\simeq 0.45$ and then rapidly compressing it.
@@
Since then, hundreds of studies have examined the structure of amorphous sphere packings with $\phi_{\rm J}\simeq\phi_{\rm MRJ}$. Denser packings have received much less attention, but the preparation-protocol dependence of $\phi_{\rm J}$ and various order parameters $\Psi$, e.g. their dependence on control variables like $\phi_{\rm init}$ and $\Gamma$, has attracted substantial interest. In particular, several studies Chaudhuri et al. (2010); Charbonneau and Morse (2021); Ozawa et al. (2017) have examined onset effects in hard sphere liquids that were first thermally equilibrated at various packing fractions $\phi_{\rm eq}$ and then jammed using “crunching” protocols that apply compression in the limit $\Gamma\to\infty$. These studies found that both $\phi_{\rm J}$ and the local structural order of jammed states (i.e., various metrics quantifying $\Psi$) increase with increasing $\phi_{\rm eq}$, provided $\phi_{\rm eq}>\phi_{\rm on}$.
@@
More specifically, Chaudhuri et al. found that the standard 50:50 1:1.4 bidisperse hard sphere model’s $\phi_{\rm J}$ increased from $0.648$ to $0.662$ as $\phi_{\rm eq}$ increased from $0.357$ to $0.584$ Chaudhuri et al. (2010). Charbonneau and Morse found that results for $\phi_{\rm J}(\phi_{\rm eq})$ can be fit by a simple functional form, and also showed that increasing $\phi_{\rm J}(\phi_{\rm eq})$ are also found in higher spatial dimensions Charbonneau and Morse (2021) despite the fact that they are absent in mean-field theories of the glass/jamming transition. Finally, Ozawa et al. exploited the SWAP Monte Carlo algorithm’s ability to equilibrate polydisperse HS liquids at larger $\phi_{\rm eq}$ Berthier et al. (2016a) to obtain much higher $\phi_{\rm J}$ Ozawa et al. (2017). For a continuously-polydisperse particle size distribution with $P(\sigma)\propto\sigma^{-3}$ for $\sigma_{\rm min}\leq\sigma\leq\sigma_{\rm max}$ and polydispersity index $\Delta=\langle\sigma^{2}\rangle/\langle\sigma\rangle^{2}-1=0.23$, they found $\phi_{\rm J}$ as high as $0.698$. These trends arise because liquids prepared at $\phi_{\rm eq}<\phi_{\rm on}$ fall out of equilibrium at $\phi_{\rm on}$ under rapid compression, but liquids prepared at $\phi_{\rm eq}>\phi_{\rm on}$ necessarily fall out of equilibrium at higher densities $\phi_{\rm g}$, and their greater structural order at these $\phi_{\rm g}$ ultimately causes them to have larger $\phi_{\rm J}$ and $\Psi$.
@@
How far can we take such effects? Refs. Chaudhuri et al. (2010); Charbonneau and Morse (2021); Ozawa et al. (2017) all employed athermal, crunching-like compression algorithms which operated in the limit $\Gamma\to\infty$. Can one obtain amorphous jammed packings with even larger $\phi_{\rm J}$ and $\Psi$ by designing a compression algorithm which effectively operates in the limit $\Gamma\to 0$? Berthier et al. recently showed that the answer is “yes” Ghimenti et al. (2024); Berthier et al. (2024). By employing an efficient event-driven Monte Carlo routine for the dynamics, and performing “cluster” SWAP (cSWAP) moves during both the initial equilibration at $\phi=\phi_{\rm eq}$ and the subsequent compression, they achieved $\phi_{\rm J}$ as large as $0.720$ for the abovementioned $P(\sigma)\propto\sigma^{-3}$ model with $\Delta=0.23$. Comparable increases in packing efficiency have been achieved for disks, using various approaches Hagh et al. (2022); bol ; Kim and Hilgenfeldt (2024). In all cases, they are achieved by employing “unphysical” moves which speed up system’s dynamics.
@@
A natural followup question that remains unanswered is: does the protocol-dependence of polydisperse spheres’ ($\phi_{\rm J},\Psi$) couple comparably strongly and nontrivially to $\Delta$? In this paper, we show that the answer to this question is “yes.” Employing standard SWAP and TDOF moves during a LS-like particle-growth process increases $\phi_{\rm J}$ by at least 6% for all $0.1\lesssim\Delta\leq 0.5$. The increases in $Z_{\rm J}$ are even more dramatic, and are strongly $\Delta$-dependent. In particular, as $\Delta$ increases, employing TDOF moves at the end of the particle-growth process increasingly allows small rattlers which would otherwise not be part of the packings’ percolating, mechanically-rigid “backbones” Moukarzel (1998) to continue growing until until they too jam, ultimately producing structures that are very dense, nearly rattler-free, and very mechanically-stable. For example, this procedure yields fractional density increases, rattler-population reductions, and bulk-modulus increases of 6.5%, 96%, and 154% for $\Delta=0.50$.
@@ II Methods
We employ the same particle size distribution
@@
used in Refs. Ozawa et al. (2017); Berthier et al. (2024); Interiano-Alberto et al. (2024) and many other recent studies of glassy dynamics Berthier et al. (2016b); Ninarello et al. (2017); Scalliet et al. (2022); Monti et al. (2022, 2023). Here $\mathcal{R}\geq 1$ is the ratio of the diameters of the largest and smallest particles. This distribution gives $\langle\sigma\rangle=2\sigma_{\rm max}/(\mathcal{R}+1)$, $\langle\sigma^{2}\rangle=2\ln(\mathcal{R})\sigma_{\rm max}^{2}/(\mathcal{R}^{2}-1)$, and
@@
Most previous studies employing this $P(\sigma)$ employed $\Delta=0.23$. Some employed smaller values, but (to the best of our knowledge) only Refs. Monti et al. (2022, 2023) have examined $\Delta>0.25$. These studies examined selected Δ=0.81, 0.92, 1.02,and $\Delta=0.81,\ 0.92,\ 1.02,\ \rm{and}\ 1.16$, and found $\phi_{\rm J}$ as high as $\simeq 0.81$ in very large systems with $\Delta\gtrsim 1$. However, they did not employ any unphysical moves during sample compression, suggesting that even higher jamming densities can be obtained when such moves are employed.
Here, to systematically characterize how employing SWAP and TDOF moves during particle growth affects ($\phi_{\rm J},\Psi$), we contrast packings generated with and without these moves for a wide range of narrowly spaced $\Delta$; specifically, $\Delta=.01i\ \forall\ i\in\{0,1,2,...,50\}$. We begin by placing $N$ hard spheres with the corresponding $P_{\mathcal{R}}(\sigma)$ randomly within cubic periodic cells while forbidding interparticle overlaps. The initial $\sigma_{\rm max}$ is chosen to be small enough that the initial packing fraction $\phi_{\rm init}<10^{-3}$. Then we employ a LS-like particle growth algorithm consisting of two stages per cycle:
@@
[…]
@@
    $N$ moves attempting to exchange the diameter of a randomly chosen particle $i$ with that of another randomly chosen particle $j$. Moves are accepted if they do not produce any interparticle overlaps.
@@
As in Ref. Hoy (2024), the attempted translations during stage (1) have maximum magnitude $0.05f$, where the move-size factor $f$ is set to $1$ at the beginning of all runs, and is multiplied by $3/4$ whenever $100$ consecutive growth cycles have passed with $\tilde{\mathcal{G}}<10^{-10}$. The LS-like algorithm terminates when $f$ drops below $2\times 10^{-9}$, which is the smallest value allowed by our double-precision numerical implementation. In the runs that do not employ SWAP, systems are considered jammed at this point.
In those that do, we reset $f$ to .01.01.01.01, and then execute a second LS-like growth process that operates like the one described above, with one critical difference: each particle is grown by a different factor, specifically $\mathcal{G}_{i}/2$, during each cycle. This procedure allows the shape of $P(\sigma)$ to vary. These “transient degree of freedom” (TDOF) moves Hagh et al. (2022); bol continue until all $N$ particles have $\mathcal{G}_{i}=0$, i.e. until no particles can grow further without producing an overlap. Systems are then considered jammed.
@@
All results presented below are averaged over at least 25 independently prepared samples.
@@ III Results
[…]
@@
Our results are entirely consistent with this qualitative picture. However, they are inconsistent with the empirical formula for continuously-polydisperse systems’ $\phi_{\rm J}$ given in Ref. Desmond and Weeks (2014):
$\phi_{\rm J}(\Delta,S)=\phi_{\rm RCP}+c_{1}\Delta+c_{2}S(\Delta)\Delta^{2},$ (5)
where $S(\Delta)$ is the skewness of systems’ particle-size distributions and $c_{1}$, $c_{2}>0$. $P_{\mathcal{R}}(\sigma)$’s skewness increases slightly faster than linearly with $\Delta(\mathcal{R})$ [Eq. 2], implying that $\phi_{\rm J}$ should increase at least cubically with $\Delta$ at large $\Delta$, but clearly this is not the case.
We do not attempt to find an alternative universal formula for $\phi_{\rm J}(\Delta)$ here – the functional form of such curves is clearly protocol-dependent – but we emphasize that our results for large $\Delta\gtrsim 0.25$ are well fit by
@@
As illustrated in Fig, 1, the fractional increases in $\phi_{\rm J}$ enabled by employing SWAP during particle growth initially grow rapidly with increasing $\Delta$, as expected. We find that the peak packing-efficiency gain ($\simeq 6.7\%$) is achieved over the range $0.13\lesssim\Delta\lesssim 0.18$. As $\Delta$ continues to increase, the gains decrease slightly, reaching a minimum of $\simeq 6.0\%$ (at $\Delta=0.39$) before gradually increasing again. The overall effect of these $\Delta$-dependent efficiency gains is to qualitatively change the shape of the $\phi_{\rm J}(\Delta)$ curve. Instead of being concave up for all $\Delta\lesssim 0.25$, it is concave down. Then, over the range $0.25\lesssim\Delta\lesssim 0.45$, it remains linear. Over this range, $\phi_{\rm J}(\Delta)$ is again fit by Eq. 6, but now with $\phi_{0}\simeq 0.669$. In other words, employing SWAP and TDOF moves enables a nearly constant packing-fraction increase $\Delta\phi=0.042$ over this range.
@@
For the well-studied case of $\Delta=0.23$, we find $\phi_{\rm J}\simeq 0.662$ ($\phi_{\rm J}\simeq 0.705$) for runs that do not (do) employ SWAP and TDOF moves. The former result is consistent with previous studies employing traditional particle-growth algorithms, low $\phi_{\rm eq}$ and moderate $\Gamma$ Berthier et al. (2016a); Ozawa et al. (2017). The latter result is consistent with the $\phi_{\rm J}$ obtained using Berthier et al.’s more-sophisticated growth algorithm with low $\phi_{\rm eq}$ Berthier et al. (2024); the higher maximum $\phi_{\rm J}\simeq 0.720$ discussed above was obtained for $\phi_{\rm eq}\simeq 0.66$.
@@
The vast majority of the increases in $\phi_{\rm J}(\Delta)$ discussed above arise from the SWAP moves. We have verified that this is so by performing independent runs (not discussed further here) that employed only SWAP moves. On the other hand, two trends shown in Fig. 1 suggest that the TDOF moves also play a critical role. First, while packing-efficiency gains decrease over the range $0.18\lesssim\Delta\lesssim 0.39$ as SWAP moves become less efficient (as expected Berthier et al. (2016a); Ninarello et al. (2017)), they increase again over the range $0.39\lesssim\Delta\lesssim 0.50$. Second, $\phi_{\rm J}(\Delta)$ increases above its linear-fit value for the largest $\Delta\gtrsim 0.45$.
@@
To better understand the above results, we examined the jammed packings’ coordination-number statistics. Our key results are shown in Figure 2. Panel (a) shows the average coordination numbers $Z_{\rm J}=N^{-1}\sum_{i=1}^{N}Z_{i}$, where particle $i$ has $Z_{i}$ contacts. In contrast to most previous studies, we do not remove rattlers (particles with $Z_{i}<4$ non ) when calculating $Z_{\rm J}$, because doing so obscures the trends we wish to highlight. Panel (b) shows the probability distributions $P(Z_{i})$, i.e. the probabilities that a given particle has $Z_{i}$ contacts. Here we define contact using the criterion $r_{ij}-\sigma_{ij}<10^{-5}$. Increasing/decreasing this contact tolerance (i.e., $10^{-5}$) shifts the $Z_{\rm J}(\Delta)$ curves up/down, lengthens/shortens the tails of the $P(Z_{i})$ distributions, and decreases/increases the rattler fractions. However, since the strengths of these shifts are only weakly preparation-protocol dependent, it does not qualitatively change the results discussed below.
Figure 2: Protocol dependence of average coordination numbers and coordination-number probability distributions for $N=1000$. In panel (b), solid and dashed curves show results obtained with and without SWAP and TDOF moves. The horizontal and vertical dashed lines respectively indicate $Z_{\rm iso}\equiv 6$, $Z_{i}=4$ and $Z_{i}=6$.

### v2

@@ Abstract
We show that for a standard continuously-polydisperse model with particle-diameter distribution $P(\sigma)\propto\sigma^{-3}$ and polydispersity index $\Delta$, employing a combination of standard SWAP moves and transient degree of freedom (TDOF) moves during a Lubachevsky-Stillinger-like particle-growth process dramatically increases the generated packings’ jamming densities $\phi_{\rm J}(\Delta)$ and coordination numbers $Z_{\rm J}(\Delta)$, for a wide range of $\Delta$. The fractional increase in $\phi_{\rm J}(\Delta)$ obtained by employing these moves first increases rapidly with $\Delta$, then plateaus at $6-7\%$ over the range $0.10\lesssim\Delta\leq 50$; the obtained $\phi_{\rm J}$ are as high as $0.747$ (for $\Delta=0.50$). These density increases are achieved without producing clearly-noticeable crystallization or increased fractionation, despite the fact that the packings are quite hyperstatic for all $\Delta>0$. SWAP and TDOF moves also reduce packings’ rattler populations by as much as 99% and increase their bulk moduli by as much as 80%.
@@ I Introduction
Issues like these led Torquato et al. to ask the question “Is random close packing well-defined?”, and demonstrate using simulations that it is not Torquato et al. (2000). They found that the varying the particle-growth rate $\Gamma$ in the Lubachevsky-Stillinger (LS) algorithm Lubachevsky et al. (1991) produces monodisperse frictionless-hard-sphere (MFHS) packings with a wide range of jamming densities $0.64\lesssim\phi_{\rm J}\lesssim 0.68$, with larger $\Gamma$ producing lower $\phi_{\rm J}$. Using these results, they proposed that jammed packings exist over a finite domain in ($\phi_{\rm J},\Psi$)-space, where $\Psi$ is some (as yet undefined) scalar structural-order parameter, and that the traditional estimate of $\phi_{\rm RCP}$ (0.637 Scott (1960)) actually corresponds to the “MRJ” packing fraction $\phi_{\rm MRJ}$, i.e, the packing fraction of maximally random jammed states that minimize $\Psi$ Torquato et al. (2000); Donev et al. (2004). In simulations of MFHS, $\phi_{\rm MRJ}$ is the jamming density obtained by starting from a state with a packing fraction below the “onset” density $\phi_{\rm on}\simeq 0.45$ and then rapidly compressing it.
@@
Since then, hundreds of studies have examined the structure of amorphous sphere packings with $\phi_{\rm J}\simeq\phi_{\rm MRJ}$. Denser packings have received much less attention, but the preparation-protocol dependence of $\phi_{\rm J}$ and $\Psi$ has attracted substantial interest. In particular, several studies Chaudhuri et al. (2010); Charbonneau and Morse (2021); Ozawa et al. (2017) have examined onset effects in hard sphere liquids that were first thermally equilibrated at various packing fractions $\phi_{\rm eq}$ and then jammed using “crunching” protocols that apply compression in the limit $\Gamma\to\infty$. These studies found that both $\phi_{\rm J}$ and various metrics quantifying $\Psi$ increase with increasing $\phi_{\rm eq}$, provided $\phi_{\rm eq}>\phi_{\rm on}$.
@@
More specifically, Chaudhuri et al. found that the standard 50:50 1:1.4 bidisperse hard sphere model’s $\phi_{\rm J}$ increased from $0.648$ to $0.662$ as $\phi_{\rm eq}$ increased from $0.357$ to $0.584$ Chaudhuri et al. (2010). Charbonneau and Morse found that results for $\phi_{\rm J}(\phi_{\rm eq})$ can be fit by a simple functional form, and also showed that increasing $\phi_{\rm J}(\phi_{\rm eq})$ are also found in higher spatial dimensions Charbonneau and Morse (2021) despite the fact that they are absent in mean-field theories of the glass/jamming transition. Finally, Ozawa et al. exploited the SWAP Monte Carlo algorithm’s ability to equilibrate polydisperse HS liquids at larger $\phi_{\rm eq}$ Berthier et al. (2016a) to obtain much higher $\phi_{\rm J}$ Ozawa et al. (2017). For a continuously-polydisperse particle size distribution with $P(\sigma)\propto\sigma^{-3}$ for $\sigma_{\rm min}\leq\sigma\leq\sigma_{\rm max}$ and polydispersity index $\Delta=\langle\sigma^{2}\rangle/\langle\sigma\rangle^{2}-1=0.23$, they found $\phi_{\rm J}$ as high as $0.698$. These trends arise because liquids prepared at $\phi_{\rm eq}<\phi_{\rm on}$ fall out of equilibrium at $\phi_{\rm on}$ under rapid compression, but liquids prepared at $\phi_{\rm eq}>\phi_{\rm on}$ necessarily fall out of equilibrium at higher densities $\phi_{\rm g}$, and their greater structural order at these $\phi_{\rm g}$ ultimately causes them to have larger $\phi_{\rm J}$ and $\Psi$ after crunching.
@@
How far can we take such effects? Refs. Chaudhuri et al. (2010); Charbonneau and Morse (2021); Ozawa et al. (2017) all employed athermal, crunching-like compression algorithms which operated in the limit $\Gamma\to\infty$. Can one obtain amorphous jammed packings with even larger $\phi_{\rm J}$ and $\Psi$ by designing a compression algorithm which effectively operates in the limit $\Gamma\to 0$? Berthier et al. recently showed that the answer is “yes” Ghimenti et al. (2024); Berthier et al. (2024). By employing an efficient event-driven Monte Carlo routine for the dynamics, and performing “cluster” SWAP (cSWAP) moves during both the initial equilibration at $\phi=\phi_{\rm eq}$ and the subsequent compression, they achieved $\phi_{\rm J}$ as large as $0.720$ for the abovementioned $P(\sigma)\propto\sigma^{-3}$ model with $\Delta=0.23$. Comparable increases in packing efficiency have been achieved for disks, using various approaches Hagh et al. (2022); bol ; Kim and Hilgenfeldt (2024). In all cases, they are achieved by employing “unphysical” moves which speed up systems’ dynamics.
@@
A natural followup question that remains unanswered is: does the protocol-dependence of polydisperse spheres’ ($\phi_{\rm J},\Psi$) couple comparably strongly and nontrivially to $\Delta$? In this paper, we show that the answer to this question is “yes.” Employing standard SWAP and TDOF moves during a LS-like particle-growth process increases $\phi_{\rm J}$ by at least 6% for all $0.1\lesssim\Delta\leq 0.5$. The increases in $Z_{\rm J}$ are even more dramatic, and are strongly $\Delta$-dependent. In particular, as $\Delta$ increases beyond $0.25$, employing TDOF moves at the end of the particle-growth process increasingly allows small rattlers which would otherwise not be part of the packings’ percolating, mechanically-rigid “backbones” Moukarzel (1998) to continue growing until they too jam, ultimately producing structures that are nearly rattler-free and very mechanically stable. For example, this procedure yields fractional density increases, rattler-population reductions, and bulk-modulus increases of 6.5%, 96%, and 80% for $\Delta=0.50$ without producing clearly-noticeable changes in commonly-investigated $\Psi$-metrics such as the pair correlation function $g(r)$ or the static structure factor $S(q)$.
The outline of the remainder of this paper is as follows. Section II describes our computational model and methods. Section III highlights several key differences between the structural and mechanical properties of packings generated with and without SWAP and TDOF moves, and analyzes how the strength and character of these differences evolves with increasing $\Delta$. In Section IV, we discuss these trends and their implications for future work. Finally, Appendices A and B provide additional details.
@@ II Methods
We employ a power-law particle size distribution
@@
that has been used in many recent studies of both glassy dynamics and jamming Ozawa et al. (2017); Berthier et al. (2024); Interiano-Alberto et al. (2024); Anzivino et al. (2023); Monti et al. (2022, 2023); Srivastava et al. (2025); Oquendo-Patino and Estrada (2020, 2022); Berthier et al. (2016b); Ninarello et al. (2017); Scalliet et al. (2022). Here $\mathcal{R}\geq 1$ is the ratio of the diameters of the largest and smallest particles. This distribution gives $\langle\sigma\rangle=2\mathcal{R}\sigma_{\rm min}/(\mathcal{R}+1)$, $\langle\sigma^{2}\rangle=2\mathcal{R}^{2}\ln(\mathcal{R})\sigma_{\rm min}^{2}/(\mathcal{R}^{2}-1)$, and
@@
Most previous studies employing this $P(\sigma)$ employed $\Delta\leq 0.23$. Ref. Anzivino et al. (2023) examined both a closely-spaced set of $0\leq\Delta\lesssim 0.65$ (a range slightly wider than that considered here) and selected larger $\Delta$. Refs. Oquendo-Patino and Estrada (2020, 2022); Monti et al. (2022, 2023); Srivastava et al. (2025) examined selected $0.20\leq\Delta\leq 1.16$, and found $\phi_{\rm J}$ as high as $\simeq 0.82$ in very large systems with $\Delta\gtrsim 1$. However, none of these studies employed unphysical moves during sample compression, suggesting that even higher jamming densities can be obtained when such moves are employed.
Here, to systematically characterize how employing SWAP and TDOF moves during particle growth affects ($\phi_{\rm J},\Psi$), we contrast packings generated with and without these moves for a wide range of narrowly spaced $\Delta$; specifically, $\Delta={\color[rgb]{0,0,1}0.01}i\ \forall\ i\in\{0,1,2,...,50\}$. We begin by placing $N$ hard spheres with the corresponding $P_{\mathcal{R}}(\sigma)$ randomly within cubic periodic cells while forbidding interparticle overlaps. The initial $\sigma_{\rm min}$ is chosen to be small enough that the initial packing fraction $\phi_{\rm init}<10^{-3}$. Then we employ a LS-like particle growth algorithm consisting of two stages per cycle:
@@
$r_{ij}$ is the center-to-center distance between particles $i$ and $j$, and $\sigma_{ij}\equiv(\sigma_{i}+\sigma_{j})/2$ is their contact distance if their diameters are $\sigma_{i}$ and $\sigma_{j}$. The minimum in Eq. 3 is taken over all neighbors of particle $i$, and the subsequent minimum defining $\tilde{\mathcal{G}}$ is taken over all $i$. These choices make the algorithm more efficient by allowing particles to grow faster when the minimal interparticle gaps are larger. Note that imposing the uniform growth rate $\tilde{\mathcal{G}}$ preserves the shape of the particle-size distribution $P_{\mathcal{R}}(\sigma)$ defined in Eq. 1. In other words, the ratio $\sigma_{\rm max}/\sigma_{\rm min}={\color[rgb]{0,0.88,0}\mathcal{R}}$ of the largest and smallest particle diameters, and indeed the ratios of all other moments of $P(\sigma)$, remain constant as $\langle\sigma\rangle=\int_{\sigma_{\rm min}}^{\sigma_{\rm max}}\sigma P(\sigma)d\sigma$ increases.
@@
    NN moves attempting to exchange the diameter of a randomly chosen particle $i$ with that of another randomly chosen particle $j$. Moves are accepted if they do not produce any interparticle overlaps.
@@
As in Ref. Hoy (2024), the attempted translations during stage (1) have maximum magnitude $0.05f$, where the move-size factor $f$ is set to $1$ at the beginning of all runs, and is multiplied by $3/4$ whenever $100$ consecutive growth cycles have passed with $\tilde{\mathcal{G}}<10^{-10}$. The LS-like algorithm terminates when $f$ drops below $2\times 10^{-9}$, which is the smallest value allowed by our double-precision numerical implementation. In the runs that do not employ TDOF moves, systems are considered jammed at this point.
In those that do, we reset $f$ to $0.01$, and then execute a second LS-like growth process that operates like the one described above, with one critical difference: each particle is grown by a different factor, specifically $\mathcal{G}_{i}/2$, during each cycle. This procedure allows the shape of $P(\sigma)$ to vary. These “transient degree of freedom” (TDOF) moves, which are similar in sprit to those performed in Refs. Hagh et al. (2022); bol , continue until all $N$ particles have $\mathcal{G}_{i}=0$, i.e. until no particles can grow further without producing an overlap. Systems are then considered jammed.
@@
All results presented below are averaged over at least 25 independently-prepared samples.
@@ III Results
III.1 Jamming densities and coordination numbers
Figure 1 shows the average jamming densities obtained with and without SWAP and TDOF moves for $N=1000$, which is the same system size employed in Ref. Berthier et al. (2024). Without these moves, small polydispersities ($\Delta\leq{\color[rgb]{0,0,1}0}.05$) produce a nearly constant $\phi_{\rm J}\simeq 0.646$, a typical value for low-dispersity frictionless spheres’ $\phi_{\rm RCP}$ Ciamarra et al. (2010). Over the range $0.05\lesssim\Delta\lesssim 0.25$, both $\phi_{\rm J}$ and $\partial\phi_{\rm J}/\partial\Delta$ gradually increase. Finally, for $\Delta\gtrsim 0.25$, $\partial\phi_{\rm J}/\partial\Delta$ remains roughly constant, and $\phi_{\rm J}(\Delta)$ increases steadily to a maximum of $\simeq 0.702$ for $\Delta=0.50$. Reasons for the increase in $\phi_{\rm J}$ with $\Delta$ have been discussed at length in Refs. Desmond and Weeks (2014); Oquendo-Patino and Estrada (2020, 2022); Monti et al. (2022, 2023); Anzivino et al. (2023); Srivastava et al. (2025). A simplistic version of the main reason is that increasing $\Delta$ allows small particles to fill in the gaps between their larger neighbors. This process, of course, does not become possible until $\Delta$ is sufficiently large.
@@ III.1 Jamming densities and coordination numbers
Our results are entirely consistent with this qualitative picture. Remarkably, they are predicted to within less than $0.3\%$ by the formula given in Eq. 25 of Ref. Anzivino et al. (2023),
$\color[rgb]{.75,0,.25}\phi_{\rm J}(\Delta)=\phi_{\rm J}(0)+\displaystyle\frac{a_{1}\Delta^{2}+a_{2}\Delta^{4}+a_{3}\Delta^{6}}{1+b_{1}\Delta^{2}+b_{2}\Delta^{4}+b_{3}\Delta^{6}},$} (5)
for all $\Delta$. Here the coefficients $\{a_{1},a_{2},a_{3},b_{1},b_{2},b_{3}\}$ represent terms that are calculated using liquid-state theory and are each functions of the monodisperse jamming density $\phi_{\rm J}(0)$ and a single fitting parameter $C_{0}$. Further details of this agreement are given in Appendix A.
Moreover, our results for $\Delta\gtrsim 0.25$ are well fit by
@@
As illustrated in Fig, 1, the fractional increases in $\phi_{\rm J}$ enabled by employing SWAP during particle growth initially grow rapidly with increasing $\Delta$, as expected. We find that the peak packing-efficiency gain ($\simeq 6.7\%$) is achieved over the range $0.13\lesssim\Delta\lesssim 0.18$. As $\Delta$ continues to increase, the gains decrease slightly, reaching a minimum of $\simeq 6.0\%$ (at $\Delta=0.39$) before gradually increasing again. The overall effect of these $\Delta$-dependent efficiency gains is to qualitatively change the shape of the $\phi_{\rm J}(\Delta)$ curve. In particular, it becomes wholly inconsistent with Eq. 5. Instead of being concave up for all $\Delta\lesssim 0.25$, it is concave down. Then, over the range $0.25\lesssim\Delta\lesssim 0.45$, it remains linear. Over this range, $\phi_{\rm J}(\Delta)$ is again fit by Eq. 6, but now with $\phi_{0}\simeq 0.669$. In other words, employing SWAP and TDOF moves enables a nearly constant packing-fraction increase $\Delta\phi=0.042$ over this range.
@@
For the well-studied case of $\Delta=0.23$, we find $\phi_{\rm J}\simeq 0.662$ ($\phi_{\rm J}\simeq 0.705$) for runs that do not (do) employ SWAP and TDOF moves. The former result is consistent with previous studies employing traditional particle-growth algorithms, low $\phi_{\rm eq}$ and moderate $\Gamma$ Berthier et al. (2016a); Ozawa et al. (2017). The latter result is consistent with the $\phi_{\rm J}$ obtained using Berthier et al.’s more-sophisticated growth algorithm with low $\phi_{\rm eq}$ Berthier et al. (2024); their higher maximum $\phi_{\rm J}\simeq 0.720$ (discussed above) was obtained for $\phi_{\rm eq}\simeq 0.66$.
@@
The vast majority of the increases in $\phi_{\rm J}(\Delta)$ discussed above arise from the SWAP moves rather than the TDOF moves. We have verified that this is so by performing independent runs (not shown in Fig. 1) that employed only SWAP moves. The additional fractional increases in $\phi_{\rm J}$ obtained during the second stage of LS-like growth (Section II) decrease from 0.07% for $\Delta=0$ to $0.01\%$ over the range $0.02\leq\Delta<0.30$. For larger $\Delta$, they increase in a concave-up fashion, reaching 0.05% by $\Delta=0.39$ and 0.3% by $\Delta=0.50$. Although these increases are small compared to those highlighted in Fig. 1’s inset, they explain two trends shown in this figure that were not discussed above: (i) while the overall packing-efficiency gains decrease over the range $0.18\lesssim\Delta\lesssim 0.39$ as SWAP moves become less efficient (as expected Berthier et al. (2016a); Ninarello et al. (2017)), they increase again over the range $0.39\lesssim\Delta\lesssim 0.50$; and (ii) $\phi_{\rm J}(\Delta)$ increases noticeably above its linear-fit value for the largest $\Delta\gtrsim 0.45$.
Figure 2: Protocol dependence of average coordination numbers and rattler fractions for $N=1000$ and $C=10^{-3}$. The horizontal dashed line indicates $Z_{\rm iso}\equiv 6$.
@@
To better understand the above results, we examined the jammed packings’ coordination-number statistics. Figure 2 shows the average coordination numbers $Z_{\rm J}=N^{-1}\sum_{i=1}^{N}Z_{i}$, where particle $i$ has $Z_{i}$ contacts. In contrast to most previous studies, we do not remove rattlers (particles with $Z_{i}<4$ non ) when calculating $Z_{\rm J}$, because doing so obscures the trends we wish to highlight. Here we define contact using the criterion employed in Ref. Ozawa et al. (2017), i.e. $r_{ij}<(1+C)\sigma_{ij}$, with $C=10^{-3}$. We chose this value because it lies within the well-known intermediate-$C$ plateau in $Z_{\rm J}(\Delta,C)$ Donev et al. (2005). Reducing the value of $C$ shifts the $Z_{\rm J}(\Delta)$ curves downwards and increases the rattler fractions, but it does not qualitatively change the results discussed below. Results for the standard definition of $Z_{\rm J}$ (i.e. the average coordination number after removing rattlers) and a further discussion of $C$-dependence are given in Appendix B.